OpenStudy (anonymous):

Which of the following equations represents a line perpendicular to y = –2x + 4? A. 2x + y = 3 B. –x + 2y = 1 C. 2x + y = 4 D. x + 2y = –4

hartnn (hartnn):

slope of y=-2x+4 is ?

OpenStudy (anonymous):

The answer is B. Use GeoGebra. Its a free download and my Algebra 2 teacher requires it. It will help a lot.

OpenStudy (anonymous):

The slope is -2

hartnn (hartnn):

slope of perpendicular line is reciprocal inverse, so, m = -1 /(-2) =1/2 gotthis ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, the equation will be of the form, y=mx+c or y=(1/2)x+c or you can determine which line has slope =1/2

OpenStudy (anonymous):

hmm... I am kind of confused but is the correct answer D?

hartnn (hartnn):

no...

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Yes the answer is B